A jar of nickels and dimes contains $3.30. There are 27 more nickels than dimes. How many of each are there?
Found 2 solutions by Paul, smoothflyin:
Answer by Paul(988)

Nickels = x+27
Dimes = x
DIme value x 100 =10
Nickel value x 100= 5
Total value x 100 =330
Equation:
5(x+27)+10(x)=330
5x+135+10x=330
15x=195
x=13

13+27=40

Hence, there are 13 dimes and 27 Nickels.
Paul.

Answer by smoothflyin(6)

To solve this first assign variables to your various terms.
Let n=nickels and let d=dimes
We know that Nickels=5 cents and dimes=10 cents so we can then say that
5n+10d=330
(5 cents times the number of nickels (n) plus 10 cents times the number of dimes (d) is equal to our total number of cents. 330 or $3.30)
we then read in our problem that There are 27 more nickels than dimes.
We can express this like so (n=d+27) or the number of nickels is equal to the number of dimes plus 27.
Now that we have a term to work with we can plug n=d+27 back into our original equattion.
5n+10d=330
5(d+27)+10d=330
5d+135+10d=330
Combine like terms
15d+135=330
Bring the "135" over the equal sign, its sign reverses and so you will subtract it from 330
15d=330-135
combine like terms
15d=195
divide through by 15
d=13
Number of dimes=13
Number of Nickles=40